find the volume of a spherical water tank with radius 8.00m
1
Expert's answer
2012-08-21T07:36:20-0400
Volume of a sphere is expressed by the formula
V=4/3*pi R^3 where R-radius,pi=3.1415926...
So we have
V=4/3*pi*8^3=2144,7 m^3